# Energy Dynamics Powered by Traction and Stress Control Formation and Motion of +1/2 Topological Defects in Epithelial Cell Monolayers

### v1

@@ Results
To study the competition between the stress and traction and the different patterns of velocity and strain rate for head-to-tail and tail-to-head defects, we recognized that computing the power density $P$ combines all of the data into a single parameter. The power density was computed according to $P=\bm{\sigma}:\dot{\bm{\varepsilon}}-\bm{t}\cdot\bm{v}/h$, with $\bm{\sigma}$ and $\dot{\bm{\varepsilon}}$ being the stress and strain rate tensors, $\bm{t}$ and $\bm{v}$ being the traction and velocity vectors, $h$ being the height of the cell layer, and the symbols :::: and $\cdot$ being the tensor and vector dot products. Our definition of $P$ is consistent with typical sign conventions in continuum mechanics, wherein a positive value of $P$ indicates dissipation of energy. Negative $P$ indicates negative dissipation, i.e., injection of energy into the system by the cells. The power density across the full cell island fluctuated notably over space, with both positive and negative power density, indicating dissipation and injection of energy, respectively (Fig. 4a). The spatial autocorrelation of power density, averaged over multiple cell islands, decayed quickly, with a spatial correlation length of 27 µm, which is approximately the width of a single cell (Fig. 4b). The small correlation distance indicates that on average, the dissipation and injection of energy is reminiscent of random noise. Interestingly, near +1/2 defects, the average power density showed clear spatial patterns. In particular, for head-to-tail defects, the power density was positive along the defect tail, indicating energy dissipation. The region of positive power density had a width of $\approx$100 µm and even existed 1 hr before defect formation (Fig. 4c,d, Supplemental Fig. S9). Tail-to-head defects exhibited negative power density, indicating energy injection, over regions of similar size (Fig. 4f,g, Supplemental Fig. S9).
[…]
@@
Following this finding, that power flows primarily along the tails of the defects, and that head-to-tail and tail-to-head defects have different signs of power flow, we hypothesized that the stress supporting elements of the cytoskeleton could differ at the tails of head-to-tail and tail-to-head defects. To test this idea, we imaged cells over time and then immediately fixed the cells and fluorescently stained actin. From the time-lapse imaging, we identified head-to-tail and tail-to-head defects, and we then collected images of actin at each of the identified defects. For head-to-tail defects, actin stress fibers were commonly present near the tail of each defect, and the stress fibers aligned with the tail. For tail-to-head defects, by contrast, actin fibers tended to align perpendicular to the tail (Fig. 4i,j Supplemental Fig. S10). To quantify these observations, we manually measured the orientation of stress fibers for cells in the tail of each defect and defined an angle $\beta$ between the orientation of the stress fibers and the defect tail (Fig. 4k). For head-to-tail defects, $\beta$ tended toward zero, indicating alignment between the cells and the tail of the defects. By contrast, for tail-to-head defects, $\beta$ showed a tendency toward 90, indicating that the stress fibers typically aligned perpendicular to the defect tail (Fig. 4l,m). Hence, cells at the tails of head-to-tail and tail-to-head defects exhibit different cytoskeletal structures that can explain the different patterns of energy transmission. For head-to-tail defects, actin stress fibers were aligned along the defect tail and potentially transmitted energy from neighboring cells into the defect. For tail-to-head defects, actin stress fibers were aligned perpendicular to the tail, and these cells acted as a source, rather than a sink, of energy.
@@ Discussion
We studied force and motion of epithelial cells near +1/2 topological defects. In contrast to prior reports, we observed that the defects moved in either the head-to-tail or tail-to-head direction, and that defects moving in both directions coexisted within the same cell monolayer. By observing that tractions at the cell-substrate interface are not solely dissipative as is often assumed, our study demonstrates how tail-to-head defects move against stress gradients, namely, they are propelled by the tractions. The resulting picture is that defect motion can be dominated either by tractions or stresses. To identify what could cause stress or traction to dominate, we quantified power density, which is a means to include both stress and traction together in a single quantity. Head-to-tail and tail-to-head defects had opposite signs of power density, indicating dissipation and injection of energy, respectively. The energy dissipated or injected near the defect was transmitted primarily along the tail of the defects. Further inspection showed that actin stress fibers were aligned with the defect tails only for head-to-tail defects, indicating that the stress fibers likely were responsible for transmission of energy from neighboring cells into the head-to-tail defects.
Our findings imply that motion of +1/2 defects is not a useful indicator of the stress state. This finding is consistent with a recent observation that the orientation of the cell body does not necessarily align with the orientation of first principal stress.26 A distinction here is that the prior study did not investigate motion of the cells or defects. Our finding that traction and motion can align also points out that traction does not act solely as a friction, as assumed by some prior models. The strong alignment between traction and motion near +1/2 defects (Fig. 3) is also surprising, because in general, it is rare for traction and velocity to align.21, 22 Our observations are with an epithelial cell type, and it will be interesting to determine whether they apply for other conditions, such as in cell types with stronger nematic order, like fibroblasts, 8 or epithelial cells with E-cadherin knocked out. 10, 24
@@
Recent studies have posed explanations for how +1/2 defects move in the tail-to-head direction, which has been thought to imply extensile behavior, even though cells are contractile. One study suggested that fluctuations in each cell’s shape can generate tail-to-head moving defects in a contractile system.34 Others showed how fluctuating polar forces, such as cell-substrate tractions, can lead to extensile behavior.35, 36 Our quantification of power provides a useful way to study the joint effect of stresses and tractions, and the data reveal notable fluctuations of power density over space. Interestingly, we found non-negligible spatial correlations in power at +1/2 defects, with the sign of power density predicting the direction of motion of the defect. Although the cause for these spatial correlations is not yet clear, spatial correlations in tractions have been observed before,23, 37 and given that the spatial correlations in traction and stress are present at $t=-1$ hr (i.e., before defect formation), it is reasonable to expect that they are important for defect formation.
@@
An important observation in this study is that patterns of strain rate, stress, traction, and power density exist in the cell monolayer even an hour before defect formation. Most notable is that the sign of strain rate and power density differ for head-to-tail and tail-to-head defects. Given these observations, it is likely that the direction of defect motion is determined before the defect forms. In the case of tail-to-head moving defects, before the defect forms, the cells apply traction in coordination with their neighbors (Fig. 3e,f), leading to collective motion and a negative strain rate (Fig. 2e,f). In turn, these cells shrink in the $x$ direction, causing them to acquire a long axis along the $y$ direction, which forms a +1/2 defect that is moving in the tail-to-head direction. In the case of the head-to-tail moving defects, locally high tensile stresses may pull on a group of cells, causing them to increase in size (positive strain rate, Fig. 2a,b). These cells resist motion with a net dissipation of energy (Fig. 4c,d), but are pulled along by energy transmitted through stress fibers along the tail of the defect (Fig. 4e, l, m).
@@
Together, our data suggest that the defects form as a result of the collective motion of a local group of cells. Nematic order would not need to exist initially—the nematic order could emerge from the collective motion. The notion that nematic order and defects result from patterns in cell motion and strain rate differs from current understanding—it is commonly assumed that a tendency for nematic order defines the local cell orientation, which in turn defines the orientation of active stresses and the resulting cell motion. Our findings reverse this picture, suggesting that cell motion causes formation of nematic order and defects. This idea is conceptually similar to recent models that showed how a balance of intercellular forces can cause cells to elongate, in turn leading to nematic order and topological defects,34, 38 though a difference is that the experiments show spatially correlated patterns of cell tractions and stresses leading to formation of topological defects. In future work, it will be interesting to study what causes these spatial patterns of cell forces.
@@ Methods
Experiments were performed with cells seeded in micropatterned islands of 1 mm diameter on polyacrylamide substrates of Young’s modulus 6 kPa having fluorescent particles localized at the top of the gel for traction force microscopy (TFM). Glass bottom dishes (Cellvis) were activated with plasma treatment and coated with 3-(Trimethoxysilyl) propyl methacrylate (MilliporeSigma). A freshly made polyacrylamide solution (5.5% w/v acrylamide, 0.2% w/v bisacrylamide, Biorad Laboratories) was mixed with 0.03% w/v fluorescent particles (diameter 0.5 µm, carboxylate modified, Life Technologies) and 50 µl of gel and bead solution was deposited on the activated glass surface. A coverslip was placed on top of the polyacrylamide solution before polymerization, and the gel was centrifuged upside down to localize the fluorescent particles to the top. The final thickness of the gels was approximately 100 µm. For micropatterning of 1 mm cell islands, polydimethysiloxane (PDMS) (Sylgard 184) was cured into a 200 µm thick sheet, and 1 mm holes were cut with a biopsy punch to create PDMS masks. The PDMS masks were sterilized with 70% ethanol and submerged in 2% Pluronic F-127 (Sigma-Aldrich) overnight at room temperature. A treated PDMS mask was placed on top of each polyacrylamide gel, and the gels were functionalized using sulfo-SANPAH (100 µg/ml, Pierce Biotechnology) to covalently crosslink rat tail collagen I (Corning). 1 ml of 0.01 mg/ml collagen was added to each gel and the gels were incubated overnight at 4C. Approximately $0.5\times 10^{6}$ cells (in 0.5 ml of medium) were put on top of each gel for 2 hr to adhere, followed by removal of the PDMS mask and replacement of the cell culture medium. The seeded cells were allowed to spread to confluence in each island for 24 hr in medium containing 2% fetal bovine serum.
@@
Time lapse imaging of cells and fluorescent particles was performed using an Eclipse Ti microscope (Nikon Instruments, Melville, NY) with a 10×10$\times$10 × numerical aperture 0.3 objective or a 20×20$\times$20 × numerical aperture 0.5 objective (Nikon) and an Orca Flash 4.0 camera (Hamamatsu, Bridgewater, NJ) controlled by Elements Ar software (Nikon). Cells were maintained at 37C and 5% CO2 using a custom-built cage incubator. Images were collected every 15 min. Immediately before imaging, some cell islands were treated with inhibitors and activators of actomyosin contraction, namely, CN03 (2 µg/mL Cytoskeleton, Inc.) and CN02 (40 ng/ml, Cytoskeleton). After imaging, cells were removed by incubating in 0.05% trypsin for 1 hr, and images of the fluorescent particles in the substrate were captured for a traction-free reference state to compute cell-substrate tractions.
@@
For imaging of stress fibers, cells in the 1 mm cell islands and their green fluorescent protein-labeled nuclei were imaged by phase contrast and fluorescence microscopy using the Eclipse Ti microscope for 8 hr and then immediately the cells were fixed in a 4% paraformaldehyde solution for 20 min and permeabilized using 0.1% Triton X. The last frame of phase contrast images was used to identify head-to-tail and tail-to-head moving +1/2 defects at the end of the 8 hr experiment (Supplemental Fig. S10a). Identified defects were plotted along with the fluorescent images of their nuclei (Supplemental Fig. S10b). Next, the fixed cells were stained for actin imaging using ActinRed 555 ReadyProbe Reagent (Invitrogen, catalog number R37112) according to manufacturer instructions. Confocal images of nuclei (Supplemental Fig. S10c) and of actin (Supplemental Fig. S10d,e) were collected on a Nikon A1R confocal microscope with a water immersion 40×40$\times$40 × numerical aperture 1.15 objective (Nikon) using NIS-Elements Ar software. The images of the nuclei were used as a guide to match the confocal images to the phase contrast images acquired during the time lapse, as described in Supplemental Fig. S10.
Data Analysis
@@
The cell velocity field $\bm{v}$ was calculated from phase contrast images of cells by using Fast Iterative Digital Image Correlation (FIDIC) 25 using $32\times 32$ pixel ($21\times 21$ µm2) subsets centered on a grid with spacing of 8 pixels (5.2 µm) and dividing by the imaging time (15 min). The vorticity was calculated from the velocity $\bm{v}$ according to $\partial v_{y}/\partial x-\partial v_{x}/\partial y$. The strain rate tensor $\dot{\bm{\varepsilon}}$ was calculated according to $\dot{\bm{\varepsilon}}=[\nabla\bm{v}+(\nabla\bm{v})^{T}]/2$, with $(\cdot)^{T}$ indicating the transpose. Numerical derivatives were computed by the central difference method. Defect identification was similar to methods reported previously 2, 6 and based on the local orientation of the cells, which was determined by analysis of phase contrast images. Phase contrast images were analyzed with the ImageJ plugin OrientationJ using a window size of 16 pix (10.4 µm) and the “Cubic Spline” option for computing the gradient. Orientation fields were plotted for manual selection of defects. Of the candidate defects identified, only +1/2 topological defects having two counter-rotating vortices on either side of the tail (as assessed by the vorticity field as in Fig. 1c,d) were used for further analysis. Defects within 100 µm of the boundaries of the cell islands were excluded from the analysis.
@@
For traction force microscopy, substrate displacements were computed by applying FIDIC to images of the fluorescent particles using $32\times 32$ pixel (21 $\times$ 21 µm2) subsets centered on a grid with spacing of 8 pixels (5.2 µm). Traction was calculated with unconstrained Fourier-transform traction microscopy accounting for substrates of finite thickness.31, 32, 33 In this manuscript, the tractions that are reported are the traction applied by the substrate to the cells. The stresses within the cell layer were calculated with monolayer stress microscopy 27, 28 (with freely available software 29) using a cell layer height of $h=5$ µm. The implementation is equivalent for elastic or viscous rheology and depends only on a single dimensionless constant, which is equal to the ratio of shear to bulk modulus (for elastic) or viscosity (for viscous) and was chosen to be 0.54 for this system. Prior work has shown dependence on this constant to be negligible.28 The use of cell islands is important, as it allows for all boundaries to be imaged such that a stress-free boundary condition can be used. Three additional Dirichlet boundary conditions are required to stabilize the computation. For a cell island for which the tractions are in a state of equilibrium, these boundary conditions do not affect the stress state, which we verified, as described previously. 29

### v2

@@ Results
The concepts of propulsion and resistance imply, respectively, injection and dissipation of energy. Following this idea, we next considered the power associated with stress and traction. To this end, we computed the power densities $P_{S}=\bm{\sigma}:\dot{\bm{\varepsilon}}$ and $P_{T}=-\bm{t}\cdot\bm{v}/h$, respectively, with $\bm{\sigma}$ and $\dot{\bm{\varepsilon}}$ being the stress and strain rate tensors, $\bm{t}$ and $\bm{v}$ being the traction and velocity vectors, $h$ being the height of the cell layer, and the symbols :: and $\cdot$ being the tensor and vector dot products. The signs used in our definitions of $P_{S}$ and $P_{T}$ are consistent with typical sign conventions in continuum mechanics, wherein a positive value indicates dissipation of energy, and negative indicates negative dissipation, i.e., injection of energy into the system by the cells. The power densities $P_{S}$ and $P_{T}$ across the full cell island fluctuated notably over space, with both positive and negative values, indicating dissipation and injection of energy, respectively (Fig. 4a,b). The spatial autocorrelation of the power densities, averaged over multiple cell islands, decayed quickly, with spatial correlation lengths of $\sim 25$ µm, which is approximately the width of a single cell (Fig. 4c). The small correlation lengths indicate that on average, the dissipation and injection of energy is reminiscent of random noise. Interestingly, near +1/2 defects, the average power densities showed clear spatial patterns. In particular, for head-to-tail defects, both $P_{S}$ and $P_{T}$ were positive along the defect tail, indicating energy dissipation. The region of positive power densities, especially $P_{T}$, had a width of $\approx$100 µm and even existed 1 hr before defect formation (Fig. 4d,e, Supplemental Fig. S9). Tail-to-head defects exhibited negative power densities, indicating energy injection, over regions of similar size (Fig. 4g,h, Supplemental Fig. S9).
Figure 4: Power flow near +1/2 defects. (a, b) Power densities associated with stress ($P_{S}$) and traction ($P_{T}$) for a representative 1 mm cell island. (c) The spatial autocorrelation of power density $C(r)$. Markers show the average over 27 different cell islands; error bars indicate the standard deviation. (d, e) Power densities $P_{S}$ and $P_{T}$ averaged over all head-to-tail defects at $t=0$ hr. White dotted squares represent the same region of interest (ROI) as in the next panel. (f) Considering a 100 µm $\times$ 100 µm square ROI centered at (25 µm, 0), the integral $h\int_{\Gamma}\bm{\sigma}\hat{\bm{n}}\cdot\bm{v}ds$ is visualized by discretizing the path $\Gamma$ into segments of length ${\Delta}s$, where ${\Delta}s$ is equal to the spatial resolution of the measured data. The colors show the quantity $E_{\gamma}=h\bm{\sigma}\hat{\bm{n}}\cdot\bm{v}{\Delta}s$, which indicates the flow of power across the boundary of the ROI. Red (positive) and green (negative) indicate power flowing into and out of the ROI, respectively. (g–i) Power densities $P_{S}$ and $P_{T}$, and power flow $E_{\gamma}$ for tail-to-head moving defects at $t=0$ hr. (j, k) Schematic side view of the power flow through the ROI, for both types of defects. Dotted lines represent the ROI. Net powers associated with stress ($E_{S}$) traction ($E_{T}$), and the net power flowing through the boundaries of the ROI ($E_{\Gamma}$) are indicated.
For these regions of energy dissipation and injection to exist, the energy must be transmitted to neighboring cells via cell-cell adhesions. Hence, we also quantified energy transmission using Clapeyron’s theorem, which relates the net energy within a region of interest to the flow of energy across the boundaries of that region. This theorem is expressed mathematically as $\int_{\Omega}P_{S}dA+\int_{\Omega}P_{T}dA=\int_{\Gamma}\bm{\sigma}\hat{\bm{n}}\cdot\bm{v}ds$, where $\Omega$ defines a region of interest with perimeter $\Gamma$, and $\hat{\bm{n}}$ is the unit outward normal vector from $\Gamma$ (see Supplemental Note 1 for a derivation). To quantify energy flow near the +1/2 defects, we defined $\Omega$ to be a $100\times 100$ µm2 region of interest (ROI) centered slightly to the right of the defect, at ($x=25$ µm, $y=0$ µm). To visualize the line integral $\int_{\Gamma}\bm{\sigma}\hat{\bm{n}}\cdot\bm{v}ds$, we discretized the path $\Gamma$ into segments of length ${\Delta}s$, where ${\Delta}s$ is equal to the spatial resolution of the measured data. We then plotted the quantity $E_{\gamma}=h\bm{\sigma}\hat{\bm{n}}\cdot\bm{v}{\Delta}s$, where the cell height $h$ is included to recover units of Watts. The magnitude of $E_{\gamma}$ was negligible on the top and bottom sides of the region of interest. Interestingly, the magnitude was largest on the right side of the defects, indicating that power flows to neighboring cells predominantly along the tail (Fig. 4f,i). For head-to-tail defects, which dissipate energy, $E_{\gamma}$ on the right side was positive, indicating a net flow of power inward and along the tail of the defects. The trend for tail-to-head defects was the opposite, with a net outward flow of power, also along the tail of the defects. To summarize these observations, we show schematically the flow of power in Fig. 4j,k. We indicate the three integrals in Clapeyron’s theorem as follows: $E_{S}=h\ \int_{\Omega}P_{S}dA$; $E_{T}=h\ \int_{\Omega}P_{T}dA$; and $E_{\Gamma}=h\ \int_{\Gamma}\bm{\sigma}\hat{\bm{n}}\cdot\bm{v}ds=\sum_{\Gamma}E_{\gamma}$. The cell layer height $h$ is included to recover units of Watts. With this new notation, Clapeyron’s theorem becomes a simple balance between three quantities, $E_{S}+E_{T}=E_{\Gamma}$. Numerical values for $E_{S}$, $E_{T}$, and $E_{\Gamma}$ are shown in Fig. 4j,k, and they balance to within a couple percent, meaning errors in the measurement are small (see Supplemental Note 2 for further quantification of noise in the measurements of power). The results show that, for head-to-tail defects, power flows across the boundaries primarily along the defect tails and is dissipated primarily at the cell-substrate interface; for tail-to-head defects, power is produced primarily at the cell-substrate interface and flows outward along the defect tails.
@@
Following the finding that power flows primarily along the tails of the defects, and that head-to-tail and tail-to-head defects have different signs of power flow, we hypothesized that the stress supporting elements of the cytoskeleton could differ at the tails of head-to-tail and tail-to-head defects. To test this idea, we imaged cells over time and then immediately fixed the cells and fluorescently stained actin. From the time-lapse imaging, we identified head-to-tail and tail-to-head defects, and we then collected images of actin at each of the identified defects. For head-to-tail defects, actin stress fibers were commonly present near the tail of each defect, and the stress fibers aligned with the tail. For tail-to-head defects, by contrast, actin fibers tended to align perpendicular to the tail (Fig. 5a,b Supplemental Fig. S10). To quantify these observations, we manually measured the orientation of stress fibers for cells in the tail of each defect and defined an angle $\beta$ between the orientation of the stress fibers and the defect tail (Fig. 5c). For head-to-tail defects, $\beta$ tended toward zero, indicating alignment between the cells and the tail of the defects. By contrast, for tail-to-head defects, $\beta$ showed a tendency toward 90, indicating that the stress fibers typically aligned perpendicular to the defect tail (Fig. 5d,e). Hence, cells at the tails of head-to-tail and tail-to-head defects exhibit different cytoskeletal structures that can explain the different patterns of energy transmission. For head-to-tail defects, actin stress fibers were aligned along the defect tail and potentially transmitted energy from neighboring cells into the defect. For tail-to-head defects, actin stress fibers were aligned perpendicular to the tail, and these cells acted as a source, rather than a sink, of energy.
Figure 5: Orientation of stress fibers near +1/2 defects. (a, b) Confocal images of stress fibers captured at the basal (bottom) of the cell layer for representative head-to-tail and tail-to-head defects. Cell boundaries were segmented using images collected at the apical (top) side of the cell layer and are shown overlaid on the images with white lines. The head and tail of each defect is marked by the dot and the line, respectively. (c) Schematic illustration showing the definition of angle $\beta$, between the orientation of stress fibers and the defect axis. (d) Histograms of angle $\beta$ for cells along the tails of head-to-tail and tail-to-head moving +1/2 defects. (e) The average of $\beta$ was computed for cells along the tail of each defect and indicated by a separate marker. Solid lines indicate means. $p=6.8\times 10^{-5}$.
@@ Discussion
We studied force and motion of epithelial cells near +1/2 topological defects. We observed that the defects moved in either the head-to-tail or tail-to-head direction, and that defects moving in both directions coexisted within the same cell monolayer. The observation of defects moving in both directions differs from prior reports in the same cell type,2, 10 and, while the cause is not yet fully understood, it may be due to the fact that we studied a system at low cell density, for which the cell monolayer is likely in the fluid state.21 By observing that tractions at the cell-substrate interface are not solely dissipative as is assumed in many theories of active nematic materials, our study demonstrates how tail-to-head defects move against stress gradients, namely, they are propelled by the tractions. The resulting picture is that defect motion can be dominated either by tractions or stresses. To study more deeply what it means for stress or traction to dominate, we quantified the power associated with stresses and traction, enabling us to study the energy dynamics. For tail-to-head and head-to-tail defects, the energy flowed in opposite directions. For tail-to-head defects, which we refer to as traction dominated, energy was injected via tractions and transmitted by stresses elsewhere in the cell layer. For head-to-tail, which we refer to as stress dominated, energy was transmitted via stresses to the defect region and dissipated via tractions at the cell–substrate interface (Fig. 4j,k). Further inspection showed that actin stress fibers were aligned with the defect tails only for head-to-tail defects, which is consistent with our observation that for these defects, the energy for motion was transmitted via stresses from neighboring cells.
A recent question of interest is how +1/2 defects can move in the tail-to-head direction, which has been thought to imply extensile behavior, even through cells are contractile. Various explanations for this unexpected behavior have been proposed. One study suggested that fluctuations in each cell’s shape can generate tail-to-head moving defects in a contractile system.41 Others showed how fluctuating polar forces, such as cell-substrate tractions, can lead to extensile behavior.22, 23 Our data showing that traction and motion can align near defects provides new insight. The data show that traction does not act solely as a friction, as assumed by some prior models. The observation of alignment between traction and motion near +1/2 defects (Fig. 3) is surprising, because traction and velocity typically do not align, except for in special situations such as the edge of an expanding cell monolayer or next to a region on the substrate to which the cells cannot adhere.40, 28 Nevertheless, traction and velocity align near tail-to-head moving defects, meaning that for these defects tractions propel the motion, and gradients in stress act as a viscous resistance.
Hence, we show that the stresses can either propel or resist defect motion, meaning that motion of +1/2 defects is not a useful indicator of the stress state. This finding is consistent with a recent observation that the orientation of the cell body does not necessarily align with the orientation of first principal stress.33 A remaining question is what defines whether the energy for defect motion comes from stresses or tractions. A hint comes from our stress fiber imaging, which shows that the alignment of stress fibers differs for head-to-tail and tail-to-head moving defects. These observations are consistent with predictions of a recent theory that proposed the existence of two distinct orientation fields in the epithelial cell layer, one for the shape of the cell body and the second for the stress-generating actomyosin fibers.42
@@
Our quantification of power provides a useful way to study the joint effect of stresses and tractions, and the data reveal notable fluctuations of power density over space. Interestingly, we found non-negligible spatial correlations in power at +1/2 defects, with the sign of power density corresponding to the direction of motion of the defect. Although the cause for these spatial correlations is not yet clear, spatial correlations in tractions have been observed before in epithelial cell monolayers29, 43 and in monolayers of bacteria near topological defects.20 Given that, in our experiments, the spatial correlations in traction and stress were present at t=1t=-1 hr (i.e., before defect formation), it is reasonable to expect that they are important for defect formation.
@@
An important observation in this study is that patterns of strain rate, stress, traction, and power density exist in the cell monolayer even an hour before defect formation. Most notable is that the signs of strain rate and power density differ for head-to-tail and tail-to-head defects. Given these observations, it is likely that the direction of defect motion is determined before the defect forms. In the case of tail-to-head moving defects, before the defect forms, the cells apply traction in coordination with their neighbors (Fig. 3e,f), leading to collective motion and a negative strain rate (Fig. 2e,f). In turn, these cells shrink in the $x$ direction, causing them to acquire a long axis along the $y$ direction, which forms a +1/2 defect that is moving in the tail-to-head direction. In the case of the head-to-tail moving defects, locally high tensile stresses may pull on a group of cells, causing them to increase in size (positive strain rate, Fig. 2a,b). These cells resist motion with a net dissipation of energy (Fig. 4d,e), but are pulled along by energy transmitted through stress fibers along the tail of the defect (Fig. 4f, and Fig. 5d,e).
@@
Together, our data suggest that the defects form as a result of the collective motion of a local group of cells. Nematic order would not need to exist initially—the nematic order could emerge from the collective motion. The notion that nematic order and defects result from patterns in cell motion and strain rate differs from current understanding—it is commonly assumed that a tendency for nematic order defines the local cell orientation, which in turn defines the orientation of active stresses and the resulting cell motion. Our findings change the focus, suggesting that cell motion may cause formation of nematic order and defects. This idea is conceptually similar to recent models that showed how a balance of intercellular forces can cause cells to elongate, in turn leading to nematic order and topological defects,41, 44 though a difference is that our experiments also show spatially correlated patterns of cell tractions and stresses leading to formation of topological defects. In future work, it will be interesting to study what causes these spatial patterns of cell forces.
@@ Methods
Experiments were performed with cells seeded in micropatterned islands of 1 mm diameter on polyacrylamide substrates of Young’s modulus 6 kPa having fluorescent particles localized at the top of the gel for traction force microscopy (TFM). Glass bottom dishes (Cellvis) were activated with plasma treatment and coated with 3-(Trimethoxysilyl) propyl methacrylate (MilliporeSigma). A freshly made polyacrylamide solution (5.5% w/v acrylamide, 0.2% w/v bisacrylamide, Biorad Laboratories) was mixed with 0.03% w/v fluorescent particles (diameter 0.5 µm, carboxylate modified, Life Technologies) and 50 µl of gel and bead solution was deposited on the activated glass surface. A coverslip was placed on top of the polyacrylamide solution before polymerization, and the gel was centrifuged upside down to localize the fluorescent particles to the top. The final thickness of the gels was approximately 100 µm. For micropatterning of 1 mm cell islands, polydimethysiloxane (PDMS) (Sylgard 184) was cured into a 200 µm thick sheet, and 1 mm holes were cut with a biopsy punch to create PDMS masks. The PDMS masks were sterilized with 70% ethanol and submerged in 2% Pluronic F-127 (Sigma-Aldrich) overnight at room temperature. A treated PDMS mask was placed on top of each polyacrylamide gel, and the gels were functionalized using sulfo-SANPAH (100 µg/ml, Pierce Biotechnology) to covalently crosslink rat tail collagen I (Corning). 1 ml of 0.1 mg/ml collagen was added to each gel and the gels were incubated overnight at 4C. Approximately $0.5\times 10^{6}$ cells (in 0.5 ml of medium) were put on top of each gel for 2 hr to adhere, followed by removal of the PDMS mask and replacement of the cell culture medium. The seeded cells were allowed to spread to confluence in each island for 24 hr in medium containing 2% fetal bovine serum. The experiments were performed at a low density of 2400 cells/mm2, which resulted in relatively long aspect ratios and greater cell motion than occurs at higher densities.21 The low density also prevented extrusion of cells out of the monolayer near +1/2 defects, which has been reported in prior studies.2, 6, 20 Extrusion events were not desirable in this study, as they could have confounding effects on the cell motion.
Microscopy
@@
Time lapse imaging of cells and fluorescent particles was performed using an Eclipse Ti microscope (Nikon Instruments, Melville, NY) with a $10\times$ numerical aperture 0.3 objective or a $20\times$ numerical aperture 0.5 objective (Nikon) and an Orca Flash 4.0 camera (Hamamatsu, Bridgewater, NJ) controlled by Elements Ar software (Nikon). Cells were maintained at 37C and 5% CO2 using a custom-built cage incubator. Images were collected every 15 min. Immediately before imaging, some cell islands were treated with inhibitors and activators of actomyosin contraction, namely, CN03 (2 µg/mL Cytoskeleton, Inc.) and CN02 (40 ng/ml, Cytoskeleton). After imaging, the cells were removed by incubating in 0.05% trypsin for 1 hr, and images of the fluorescent particles in the substrate were captured for a traction-free reference state to compute cell-substrate tractions.
@@
For imaging of stress fibers, cells in the 1 mm cell islands and their green fluorescent protein-labeled nuclei were imaged by phase contrast and fluorescence microscopy using the Eclipse Ti microscope for 8 hr and then immediately the cells were fixed in a 4% paraformaldehyde solution for 20 min and permeabilized using 0.1% Triton X-100. The last frame of phase contrast images was used to identify head-to-tail and tail-to-head moving +1/2 defects at the end of the 8 hr experiment (Supplemental Fig. S10a). Identified defects were plotted along with the fluorescent images of their nuclei (Supplemental Fig. S10b). Next, the fixed cells were stained for actin imaging using ActinRed 555 ReadyProbe Reagent (Invitrogen, catalog number R37112) according to manufacturer instructions. Confocal images of nuclei (Supplemental Fig. S10c) and of actin (Supplemental Fig. S10d,e) were collected on a Nikon A1R confocal microscope with a water immersion $40\times$ numerical aperture 1.15 objective (Nikon) using NIS-Elements Ar software. The images of the nuclei were used as a guide to match the confocal images to the phase contrast images acquired during the time lapse, as described in Supplemental Fig. S10.
Quantification of Cell Velocity and Orientation
@@
The cell velocity field $\bm{v}$ was calculated from phase contrast images of cells by using Fast Iterative Digital Image Correlation (FIDIC) 32 using $32\times 32$ pixel ($21\times 21$ µm2) subsets centered on a grid with spacing of 8 pixels (5.2 µm) and dividing by the imaging time (15 min). The vorticity was calculated from the velocity $\bm{v}$ according to $\partial v_{y}/\partial x-\partial v_{x}/\partial y$. The strain rate tensor $\dot{\bm{\varepsilon}}$ was calculated according to $\dot{\bm{\varepsilon}}=[\nabla\bm{v}+(\nabla\bm{v})^{T}]/2$, with $(\cdot)^{T}$ indicating the transpose. Numerical derivatives were computed by the central difference method. Defect identification was similar to methods reported previously 2, 6 and based on the local orientation of the cells, which was determined by analysis of phase contrast images. Phase contrast images were analyzed with the ImageJ plugin OrientationJ using a window size of 16 pix (10.4 µm) and the “Cubic Spline” option for computing the gradient. Orientation fields were plotted for manual selection of defects. To rule out that candidate defects identified were due to errors in quantifying the cell orientations, we used the vorticity as a verification criterion, and kept only +1/2 topological defects having two counter-rotating vortices on either side of the tail (as in Fig. 1c,d) for further analysis. Given that the cells tend to align with boundaries of the cell island and the cell motion near boundaries is restricted to the azimuthal direction, defects within 100 µm of the boundaries of the cell islands were excluded from the analysis.
Quantification of Tractions and Stresses
@@
For traction force microscopy, substrate displacements were computed by applying FIDIC to images of the fluorescent particles using $32\times 32$ pixel (21 $\times$ 21 µm2) subsets centered on a grid with spacing of 8 pixels (5.2 µm). Traction was calculated with unconstrained Fourier-transform traction microscopy accounting for substrates of finite thickness.38, 39, 40 In this manuscript, the tractions that are reported are the traction applied by the substrate to the cells. The stresses within the cell layer were calculated with monolayer stress microscopy 34, 35 (with freely available software 36) using a cell layer height of $h=5$ µm. Briefly, monolayer stress microscopy computes the three components of the in-plane stress tensor from the traction data. Two of the three equations required are force equilibrium in the two in-plane directions, which make no assumptions about the rheology of the cell layer or the stress state. The third equation comes from the compatibility equation of continuum mechanics, which enforces smoothness over space. Applying this third equation assumes a linear, passive relationship between stress and strain rate. The implementation is equivalent for elastic or viscous rheology and depends only on a single dimensionless constant, which is equal to the ratio of shear to bulk modulus (for elastic) or viscosity (for viscous) and was chosen to be 0.54 for this system. Prior work has shown dependence on this constant to be negligible.35 Given that the active component of the stresses is not accounted for in the compatibility equation used in monolayer stress microscopy, it is important to consider potential errors. A theoretical model for collective migration studied these errors,45 with results suggesting that if the magnitude of traction is sufficiently large (e.g., at least 1/3 of the magnitude of stress), errors in the stresses calculated by monolayer stress microscopy become negligible. As shown in Supplementary Fig. S1, this requirement is met. Our experiments used cell islands, which is important for monolayer stress microscopy, as it allows for all boundaries to be imaged such that a stress-free boundary condition can be used. Three additional Dirichlet boundary conditions are required to stabilize the computation. For a cell island for which the tractions are in a state of equilibrium, these boundary conditions do not affect the stress state, which we verified, as described previously. 36
Spatial Autocorrelation